OpenStudy (anonymous):

Illustrate the use of the quadratic formula to solve the same equation: x² + 6x + 9 = 25

OpenStudy (anonymous):

\[\huge x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (kkutie7):

x^2+6x-16 is the same as what you have.

OpenStudy (anonymous):

Quadratic formula, plug and chug.

OpenStudy (kkutie7):

A=1 B=6 and C=-16. plug into what @iambatman gave you.

OpenStudy (kkutie7):

x= 2, -8...... so you can check your work.